Please show me how to find the x-intercept and y-intercept of the line 2x+3y=-18
Answer by PRMath(133)

x-intercept and y-intercept of the line 2x+3y=-18

Here are two facts to keep in mind:

When you want to find the X intercept SOLVE for X by making y = 0
When you want to find the Y intercept SOLVE for Y by making x = 0.

It's that simple. Just plug in 0... like this:

Let's find the X intercept first. We'll SOLVE FOR X by making y = 0.

2x+3y=-18 (original equation)
2x + 3(0) = -18 (plug in 0 for the y variable)
2x = -18 (3 times 0 equals 0, so we just wipe it out of the equation).
x = -9 (Divide both sides by 2 to isolate the x. So, -18 divided by 2 is -9)

Now we know, when y =0, then x = -9. So our X intercept is: (-9, 0)


Now let's find the Y intercept. We'll SOLVE FOR Y by making x = 0.
2x+3y=-18 (original equation)
2(0) + 3y = -18 (plug in 0 for the X variable)
3y = -18 (2 times 0 equals 0, so we just wipe it out of the equation).
y = -6 (Divide both sides by 3 to isolate the y. So.. -18 divided by 3 is -6)

Now we know, when x = 0, then y = -6. So our Y intercept is: (0, -6)

That's all there is to it. I hope this was helpful. :-)